# Using dust to constrain dark matter models

### v1

@@ 2.3 Baryonic model
We make use of the baryonic prescriptions that come as part of the public release of gadget-4. These include a simplified version of the Springel & Hernquist (2003) model for star formation and supernova, and the Katz et al. (1996) model for radiative cooling. The star formation model is a multi-phase model with prescriptions for the separate treatment of hot and cold gas, forming stars from cold gas after reaching a pressure threshold, and adopts a Salpeter (1955) initial mass function (IMF). Supernovae are assumed to happen within a single timestep in the simulation, where some fraction of the cold gas designated for star formation is immediately transferred to the hot gas phase. The cooling module treats hydrogen and helium atomic cooling in collisional equilibrium, with the abundances of the ionisation states calculated using an external ultraviolet background field, following a modified Haardt & Madau (1996) spectrum (Davé et al., 1999). We adopt this simple baryonic model for straightforward comparisons between our simulations. The baryonic model doesn’t include stellar winds or supernova-driven outflows, which is simplistic compared to state-of-the-art models (e.g. Crain et al., 2015; Pillepich et al., 2018; Davé et al., 2019), but more complex models require more complex analysis beyond the scope of this analysis.
We run two simulations for each dark matter type, with different supernovae efficiency parameters, four in total. This is to determine which parameters create the most degenerate solution in the satellite populations when compared with local satellite observations. These are efficiencies of 10%, the fiducial value quoted in Springel & Hernquist (2003) for a Salpeter IMF, and 5%, knowing that a reduced supernovae efficiency is required in a warm dark matter simulation to produce similar numbers of low mass galaxies as a cold dark matter simulation. As highlighted in Section 1, the efficiencies here refer to the parameter used in Springel & Hernquist (2003), which is the mass fraction of new stars that become supernova within one simulation timestep.
@@ 2.5 Gini coefficients
The Gini coefficient is used to characterise the spatial distribution of dust produced in each simulation. Gini coefficients typically measure uniformity in a sample. When used in processes such as galaxy imaging, the scores typically represent photon flux across pixels in a CCD, where a value of 0 would mean perfect equality in the flux across all pixels, and 1 indicates all flux exists in a single pixel (Abraham et al., 2003; Lotz et al., 2004; Zamojski et al., 2007).
In our analysis, we use the Gini coefficient to describe the concentration of dust in the halo. We create a 2D projection map of the dust mass in the halo, with a bin width of 2.5ckpc, and use the total mass in each bin to calculate the Gini score, G:
where is the value in the nth cell, and is the average over cells. The dust mass is calculated out to a radius of ckpc, which equates roughly twice R at ; however, bins outside of R or within a circle with a 15ckpc radius at the centre of the halo are masked to exclude them from the Gini calculation. The former avoids counting bins beyond the halo with no dust, which would lower the Gini score, and the latter avoids the highly concentrated central galaxy, which saturates any signal. To avoid projection effects, where satellites might lie directly behind or in front of the central galaxy, we rotate the dust projection maps around each axis. We use 25 viewing angles between 0 and 2 in the x-y, x-z, and y-z axes, then take the median of the 75 individual Gini score calculations. The masking is applied in every projection at all viewing angles. This method is explored further in Appendix B.
@@ 3.1 Milky Way reproduction
Before analysing the simulations, we convert their comoving units to physical units and remove the dependence to compare with each other and later observational data. We then verify that our simulations all produce similar halo properties at . Each pair of dark matter simulations produces very similar Milky Way-like galaxies, with percent level differences in radius, total mass, and dark matter mass, which are summarised in Table 2. There are minor differences in the gas and stellar mass, which is to be expected from the differing supernova efficiencies. Comparing the two dark matter models against each other, we see that cold dark matter produces slightly larger haloes on average than warm dark matter, approximately 10% larger in total mass and 3% larger radii. Minor differences between the haloes are to be expected, as the different dark matter models natively produce different formation histories, and the supernova efficiencies give each galaxy a unique evolution. However, these differences between the simulations are within the expected numerical variance.
Table 2 also shows the observed values for selected Milky Way properties, taken from Bland-Hawthorn & Gerhard (2016), against which we can compare the four simulations. Our resimulation target selection process ensures that we recover a total halo mass in the range of 1-2. Our stellar masses are generally consistent, though a little larger than the observations, which tend to fall between 4-7 M⊙ (see also Kafle et al., 2014; Licquia & Newman, 2015; McMillan, 2017). Star formation rates also agree with the observations, which are between 1 to 3 M⊙ yr-1 (Chomiuk & Povich, 2011; Licquia & Newman, 2015; Elia et al., 2022). See Bland-Hawthorn & Gerhard (2016) for a review of the Milky Way properties.
@@ 3.2 Calibrating the simulations
The most significant point of difference between the dark matter models is the satellite galaxies, which we now compare in our simulations. We plot the cumulative satellite mass function for the Milky Way analogues in Figure 1, with the ten brightest Milky Way satellites overplotted in red triangles. We use observed satellite masses from Table 1 of Lovell & Zavala (2023) and references therein. For this work, we define a satellite galaxy in the simulation as a subhalo with at least one stellar particle and at least 100 total particles, corresponding to a minimum stellar mass of M⊙ and a minimum total mass of M⊙. The minimum total particle number limit is implemented to try to avoid artificial haloes known to plague warm dark matter simulations (Wang & White, 2007; Power et al., 2016). In this setup, these artificial haloes will dominate the mass function at M⊙ (Wang & White, 2007; Lovell et al., 2014), which is well below our halo mass resolution limit. While we include anything with a singular star particle, most satellites contain 10 star particles and generally several hundred gas and dark matter particles. Contrary to expectations, the warm dark matter simulation with 5% supernova efficiency produces the most satellite galaxies of all the simulations. We expect this is a unique case where the satellites have merged at different times rather than a trend in the simulation. We discuss this further in Section 4.2. All the simulations show good agreement with the observed number of satellites, but with so few, it’s difficult to distinguish the effect of the combination of dark matter models and feedback efficiencies.
To better highlight the satellite population degeneracy between simulations, in Figure 2, we present the entire population of satellite galaxies. This figure shows a similar cumulative mass function, but now for all satellites in the box rather than only those around the Milky Way analogue. In order to be classified as a satellite in this context, the FOF group must contain no low-resolution dark matter particles. Additionally, the subhalo must contain at least one star particle and must not be rank 0, the subfind convention for identifying the host halo. We see the expected divergence at lower masses, with both warm dark matter simulations producing fewer satellites overall than their cold dark matter counterparts. The different supernova efficiencies also show the expected behaviour, with higher rates producing fewer satellites. We find a relatively consistent match between the cold dark matter simulation with a 10% supernova efficiency rate and the warm dark matter simulation with a 5% rate. We designate these two simulations as our fiducial models for our analysis below.
@@ 3.3 Dust comparison across the simulations
Table 3 shows the total dust mass across all simulations, within R and the central 15kpc, using the dust-to-metals prescription described in Section 2.4. Crucially, the difference between the fiducial cold (10%) and warm (5%) dark matter models (the two best matching configurations) is significant, with times the total amount of dust in the cold dark matter simulation. An increase in supernovae efficiency increases the amount of dust in both dark matter models, though less so for warm dark matter. The lower supernova efficiency effectively locks away more gas in stars, reducing the availability of gas for star-forming episodes that drive increases in metallicity. The concentration, or the fraction of dust within 15kpc of the halo centre, ranges from 75% in the fiducial warm dark matter simulation to up to 87% in the fiducial cold dark matter simulation. This presents the possibility that direct observations of the dust inside a galaxy can be used as a metric to differentiate the nature of the dark matter species, in terms of aggregated dust masses.
@@
To explore dust outside of the galaxy (but inside the halo), Figures 3(a) and 3(b) plot the 2D projections of its distribution out to the halo radius for our fiducial cold and warm dark matter simulations. Beyond the central galaxy, the remaining dust tends to be concentrated in the satellites, with lower densities in the diffuse gas. The differential map between the two simulations is shown in Figure 3(c). The colour bar shows the percentage difference between the two, with blue indicating higher dust mass in the cold dark matter simulation and red indicating higher dust mass in the warm dark matter simulation. This comparison shows the same behaviour as that of the total mass, that generally, there is more dust in the cold dark matter simulation spread throughout the halo. Besides this general interpretation, the differential map has several key features. The most obvious is that the difference in satellite positions creates stark differences in the differential dust map, even if their positions are only marginally offset. The other interesting feature is the difference in dust distribution around the central galaxy. The cold dark matter simulation has a much more compact dust disk than the warm dark matter simulation, which is primarily driven by the difference in supernova efficiencies. When we switch the efficiencies, i.e. cold dark matter with a 5% supernova efficiency and warm dark matter with a 10% efficiency, we see the opposite trend, where the cold dark matter has a much more extended dust disk. This switch doesn’t significantly affect the dust outside of the central galaxy.
@@
To better understand the difference in dust mass between the simulations, we plot the dust surface density as a function of the radius at in Figure 4. We over-plot the scaling relation from Ménard et al. (2010) with a scaling amplitude =0.03 for the dust-to-metals ratio, chosen to match the densities of the simulations at radii kpc, as well as a red arrow the point where gravity becomes Newtonian, and a dashed line at the masking radius. We also plot the radial dust surface density of M31, the closest Milky Way analogue to the Milky Way Galaxy. The surface density is taken from Draine et al. (2014), where they find a total estimated dust mass of M⊙. Our results generally agree with the M31 observations in amplitude and shape at intermediate radii, between 10-20kpc, although inner regions have higher densities than M31. The high-density central regions are potentially inflated by limitations of our baryonic model; future refinements could add stellar winds or AGN feedback to help drive gas into the circumgalactic medium and lower the dust content in the central galaxy. The cold dark matter simulation with 10% supernova efficiency is additionally inflated by a recent close pass by a merging satellite, which happens at larger radii for the other cold dark matter simulation and not at all in the warm dark matter simulations. All simulations follow the Ménard et al. (2010) scaling relation at radii greater than 20kpc with little deviation. Outside of the central galaxy, all simulations have similar surface density profiles.
@@ 3.4 Gini results
In Section 2.5, we introduced the Gini score, which we now apply to our projected dust surface maps to further discriminate between their distributions and how this changes with time. The Gini score allows us to condense the information in maps like those in Figures 3(a) and 3(b), into a single number. This is calculated for all halo rotations (see Section 2.5) between and , and is plotted in the top panel of Figure 5. The shaded regions in the plot show 1 bootstrap errors for the medians at each redshift. The relative difference in Gini score with respect to the 10% cold dark matter simulation is shown in the bottom panel of Figure 5, with the same errors as above.
At higher redshift, cold and warm dark matter simulations share similar Gini scores and evolutions, but this begins to diverge around . Both warm dark matter simulations have lower concentrations of dust and hence a lower score than both cold dark matter simulations, and the gap between them grows with time. Supernova feedback efficiency doesn’t appear to significantly impact the Gini score for a given dark matter type. The stochastic nature of the overall score is primarily governed by the merger history; for a universe such as warm dark matter with fewer low mass haloes and fewer merger events, we expect reduced Gini scores compared to a purely cold dark matter universe. This is a promising result as a proof of concept. Future work will focus on a wider range of Milky Way-like haloes and galaxy histories to explore the significance of the variance of such statistics.
@@ 4.1 Simulations
Our simulations and subsequent dust post-processing show that we have potentially uncovered a new observational probe to test the nature of dark matter. Through the use of varying supernova efficiencies as our lever, we created simulations from different underlying dark matter types that produce the same “observable” satellite populations. These different supernova efficiencies lead to different dust properties in each halo. We find that the total dust mass produced by the fiducial cold dark matter simulation, with a 10% supernova efficiency, is 4 times greater than in the fiducial warm dark matter simulation. We also presented a preliminary exploration of how dust is spatially distributed within the halo using the Gini statistic, which shows significant and persistent differences between dark matter models. The difference between the two suggests that dust can be used as an observational probe of the underlying dark matter. In reality, galactic dust masses are influenced by many factors, such as star formation histories (Calzetti et al., 2000; Clemens et al., 2013), mergers (Buck et al., 2023), and environments (Chelouche et al., 2007; Vogelsberger et al., 2019; Clark et al., 2023), so more work must be done before we draw conclusions about any observationally favoured model.
@@ 4.2 Milky Way satellites
There are many different sources for the mass of Milky Way satellites, which introduces some uncertainty into the correct values against which to calibrate the simulations. We use the masses published in Lovell & Zavala (2023) for all satellites included in the comparison in Figure 1. This maintains a single source of data for simplicity and transparency, but does not explore the uncertainty of masses for these satellites. Many stellar masses are taken from McConnachie (2012), which has minor differences with other published masses for these satellites (Woo et al., 2008; Fattahi et al., 2018), though all are quite consistent for the larger satellites. Additionally, there is some scatter in the dynamical masses compared with Strigari et al. (2008) and Wolf et al. (2010), but once again, being the largest Milky Way satellites, the scatter in these measurements is relatively small.
Counterintuitively, Figure 1 shows that the fiducial warm dark matter simulation produces more low-mass satellite galaxies than any other simulation, as mentioned in Section 3.2. This is rather unexpected, given that warm dark matter is generally implemented to decrease satellite numbers. The apparent overproduction in the simulation 222Although we note that both warm dark matter simulations produce fewer total subhaloes than their cold dark matter counterparts. could be due to a number of factors, first among them the artificial fragmentation common in warm dark matter simulations (Wang & White, 2007). We have explored removing such artificial halos by implementing a minimum 100-particle limit. Our testing suggests that artificial halos are not the cause of the excess. More work is needed to identify spurious haloes, e.g. identifying pancake haloes in the initial conditions (Lovell et al., 2014), but this is beyond the scope of this paper. Alternatively, Oman et al. (2024) found that warm dark matter may boost the number of low-mass galaxies in 21-cm hydrogen observations. They state this may be because galaxies tend to form later in warm dark matter simulations, and the subsequent delay in star formation helps them retain more gas than their cold dark matter counterparts. Additionally, the higher concentration of cold dark matter haloes may drive faster merging times, causing an excess of warm dark matter satellites that have yet to fully merge by the same point in time. This may be true, at least for our Milky Way analogue. We defer such an investigation to later work, conscious that the most stringent limits on warm dark matter particle masses are based on the observation of satellites below the resolution of these simulations (Banik et al., 2021; Nadler et al., 2021a; Dekker et al., 2022). However, the possibility of a warm dark matter universe potentially producing more satellite galaxies in some haloes prompts further investigation into the variety of pathways for satellite accretion in a warm dark matter universe and how these might compare with our own Milky Way satellites.

### v2

@@ 2.3 Baryonic model
We make use of the baryonic prescriptions that come as part of the public release of gadget-4. These include a simplified version of the Springel & Hernquist (2003) model for star formation and supernova, and the Katz et al. (1996) model for radiative cooling. The star formation model is a multiphase model with prescriptions for the separate treatment of hot and cold gas, forming stars from cold gas after reaching a pressure threshold, and adopts a Salpeter (1955) initial mass function (IMF). Supernovae are assumed to happen within a single time-step in the simulation, where some fraction of the cold gas designated for star formation is immediately transferred to the hot gas phase. The cooling module treats hydrogen and helium atomic cooling in collisional equilibrium, with the abundances of the ionization states calculated using an external ultraviolet background field, following a modified Haardt & Madau (1996) spectrum (Davé et al., 1999). We adopt this simple baryonic model for straightforward comparisons between our simulations. The baryonic model doesn’t include stellar winds or supernova-driven outflows, which is simplistic compared to state-of-the-art models (e.g. Crain et al., 2015; Pillepich et al., 2018; Davé et al., 2019), but more complex models require more complex analysis beyond the scope of this analysis.
We run two simulations for each dark matter type, with different supernovae efficiency parameters, four in total. This is to determine which parameters create the most degenerate solution in the satellite populations when compared with local satellite observations. These are efficiencies of 10 per cent the fiducial value quoted in Springel & Hernquist (2003) for a Salpeter IMF, and 5 per cent, knowing that a reduced supernovae efficiency is required in a WDM simulation to produce similar numbers of low mass galaxies as a CDM simulation. As highlighted in Section 1, the efficiencies here refer to the parameter used in Springel & Hernquist (2003), which is the mass fraction of new stars that become supernova within one simulation time-step.
@@ 2.5 Gini coefficients
The Gini coefficient is used to characterize the spatial distribution of dust produced in each simulation. Gini coefficients typically measure uniformity in a sample. When used in processes such as galaxy imaging, the scores typically represent photon flux across pixels in a CCD, where a value of 0 would mean perfect equality in the flux across all pixels, and 1 indicates all flux exists in a single pixel (Abraham et al., 2003; Lotz et al., 2004; Zamojski et al., 2007).
In our analysis, we use the Gini coefficient to describe the concentration of dust in the halo. We create a 2D projection map of the dust mass in the halo, with a bin width of ckpc, and use the total mass in each bin to calculate the Gini score, :
where is the value in the th cell, and is the average over cells. The dust mass is calculated out to a radius of ckpc, which equates roughly twice at ; however, bins outside of or within a circle with a 15ckpc radius at the centre of the halo are masked to exclude them from the Gini calculation. The former avoids counting bins beyond the halo with no dust, which would lower the Gini score, and the latter avoids the highly concentrated central galaxy, which saturates any signal. To avoid projection effects, where satellites might lie directly behind or in front of the central galaxy, we rotate the dust projection maps around each axis. We use 25 viewing angles between 0 and 2 in the , , and axes, then take the median of the 75 individual Gini score calculations. The masking is applied in every projection at all viewing angles. This method is explored further in Appendix B.
@@ 3.1 Milky Way reproduction
Before analysing the simulations, we convert their comoving units to physical units and remove the dependence to compare with each other and later observational data. We then verify that our simulations all produce similar halo properties at . Each pair of dark matter simulations produces very similar Milky Way-like galaxies, with per cent level differences in radius, total mass, and dark matter mass, which are summarized in Table 2. There are minor differences in the gas and stellar mass, which is to be expected from the differing supernova efficiencies. Comparing the two dark matter models against each other, we see that CDM produces slightly larger haloes on average than WDM, approximately 10 per cent larger in total mass and 3 per cent larger radii. Minor differences between the haloes are to be expected, as the different dark matter models natively produce different formation histories, and the supernova efficiencies give each galaxy a unique evolution. However, these differences between the simulations are within the expected numerical variance.
Table 2 also shows the observed values for selected Milky Way properties, taken from Bland-Hawthorn & Gerhard (2016), against which we can compare the four simulations. Our resimulation target selection process ensures that we recover a total halo mass in the range of 1-2. Our stellar masses are generally consistent, though a little larger than the observations, which tend to fall between 4 and 7 M⊙ (see also Kafle et al., 2014; Licquia & Newman, 2015; McMillan, 2017). Star formation rates also agree with the observations, which are between 1 and 3 M⊙ yr-1 (Chomiuk & Povich, 2011; Licquia & Newman, 2015; Elia et al., 2022). See Bland-Hawthorn & Gerhard (2016) for a review of the Milky Way properties.
@@ 3.2 Calibrating the simulations
The most significant point of difference between the dark matter models is the satellite galaxies, which we now compare in our simulations. We plot the cumulative satellite mass function for the Milky Way analogues in Fig. 1, with the ten brightest Milky Way satellites overplotted in red triangles. We use observed satellite masses from Table 1 of Lovell & Zavala (2023) and references therein. For this work, we define a satellite galaxy in the simulation as a subhalo with at least one stellar particle and at least 100 total particles, corresponding to a minimum stellar mass of M⊙ and a minimum total mass of M⊙. The minimum total particle number limit is implemented to try to avoid artificial haloes known to plague WDM simulations (Wang & White, 2007; Power et al., 2016). In this setup, these artificial haloes will dominate the mass function at M⊙ (Wang & White, 2007; Lovell et al., 2014), which is below our satellite mass resolution limit. While we include anything with a singular star particle, most satellites contain 10 star particles and generally several hundred gas and dark matter particles. Contrary to expectations, the WDM simulation with 5 per cent supernova efficiency produces the most satellite galaxies of all the simulations. We expect this is a unique case where the satellites have merged at different times rather than a trend in the simulation. We discuss this further in Section 4.2. All the simulations show good agreement with the observed number of satellites, but with so few, it’s difficult to distinguish the effect of the combination of dark matter models and feedback efficiencies.
To better highlight the satellite population degeneracy between simulations, in Fig. 2, we present the entire population of satellite galaxies. This figure shows a similar cumulative mass function, but now for all satellites in the box rather than only those around the Milky Way analogue. In order to be classified as a satellite in this context, the FOF group must contain no low-resolution dark matter particles. Additionally, the subhalo must contain at least one star particle and must not be rank 0, the subfind convention for identifying the host halo. We see the expected divergence at lower masses, with both WDM simulations producing fewer satellites overall than their CDM counterparts. The different supernova efficiencies also show the expected behaviour, with higher rates producing fewer satellites. We find a relatively consistent match between the CDM simulation with a 10 per cent supernova efficiency rate and the WDM simulation with a 5 per cent rate. We designate these two simulations as our fiducial models for our analysis below.
@@ 3.3 Dust comparison across the simulations
Table 3 shows the total dust mass across all simulations, within and the central 15kpc, using the dust-to-metals prescription described in Section 2.4. Crucially, the difference between the fiducial cold (10 per cent) and warm (5 per cent) dark matter models (the two best matching configurations) is significant, with times the total amount of dust in the CDM simulation. An increase in supernovae efficiency increases the amount of dust in both dark matter models, though less so for WDM. The lower supernova efficiency effectively locks away more gas in stars, reducing the availability of gas for star-forming episodes that drive increases in metallicity. The concentration, or the fraction of dust within 15kpc of the halo centre, ranges from 75 per cent in the fiducial WDM simulation to up to 87 per cent in the fiducial CDM simulation. This presents the possibility that direct observations of the dust inside a galaxy can be used as a metric to differentiate the nature of the dark matter species, in terms of aggregated dust masses.
@@
To explore dust outside of the galaxy (but inside the halo), Figs 3(a) and 3(b) plot the 2D projections of its distribution out to the halo radius for our fiducial cold and warm dark matter simulations. Beyond the central galaxy, the remaining dust tends to be concentrated in the satellites, with lower densities in the diffuse gas. The differential map between the two simulations is shown in Fig. 3(c). The colour bar shows the percentage difference between the two, with blue indicating higher dust mass in the CDM simulation and red indicating higher dust mass in the WDM simulation. This comparison shows the same behaviour as that of the total mass, that generally, there is more dust in the CDM simulation spread throughout the halo. Besides this general interpretation, the differential map has several key features. The most obvious is that the difference in satellite positions creates stark differences in the differential dust map, even if their positions are only marginally offset. The other interesting feature is the difference in dust distribution around the central galaxy. The CDM simulation has a much more compact dust disc than the WDM simulation, which is primarily driven by the difference in supernova efficiencies. When we switch the efficiencies, i.e. CDM with a 5 per cent supernova efficiency and WDM with a 10 per cent efficiency, we see the opposite trend, where the CDM has a much more extended dust disc. This switch does not significantly affect the dust outside of the central galaxy.
@@
To better understand the difference in dust mass between the simulations, we plot the dust surface density as a function of the radius at in Fig. 4. We overplot the scaling relation from Ménard et al. (2010) with a scaling amplitude =0.03 for the dust-to-metals ratio, chosen to match the densities of the simulations at radii kpc, as well as a red arrow the point where gravity becomes Newtonian, and a dashed line at the masking radius. We also plot the radial dust surface density of M31, the closest Milky Way analogue to the Milky Way Galaxy. The surface density is taken from Draine et al. (2014), where they find a total estimated dust mass of M⊙. Our results generally agree with the M31 observations in amplitude and shape at intermediate radii, between 10 and 20kpc, although inner regions have higher densities than M31. The high-density central regions are potentially inflated by limitations of our baryonic model; future refinements could add stellar winds or AGN feedback to help drive gas into the circumgalactic medium and lower the dust content in the central galaxy. The CDM simulation with 10 per cent supernova efficiency is additionally inflated by a recent close pass by a merging satellite, which happens at larger radii for the other CDM simulation and not at all in the WDM simulations. All simulations follow the Ménard et al. (2010) scaling relation at radii greater than 20kpc with little deviation. Outside of the central galaxy, all simulations have similar surface density profiles.
@@ 3.4 Gini results
In Section 2.5, we introduced the Gini score, which we now apply to our projected dust surface maps to further discriminate between their distributions and how this changes with time. The Gini score allows us to condense the information in maps like those in Figures 3(a) and 3(b), into a single number. This is calculated for all halo rotations (see Section 2.5) between and , and is plotted in the top panel of Fig. 5. The shaded regions in the plot show 1 bootstrap errors for the medians at each redshift. The relative difference in Gini score with respect to the 10 per cent CDM simulation is shown in the bottom panel of Fig. 5, with the same errors as above.
At higher redshift, cold and warm dark matter simulations share similar Gini scores and evolutions, but this begins to diverge around . Both WDM simulations have lower concentrations of dust and hence a lower score than both CDM simulations, and the gap between them grows with time. Supernova feedback efficiency doesn’t appear to significantly impact the Gini score for a given dark matter type. The stochastic nature of the overall score is primarily governed by the merger history; for a universe such as WDM with fewer low mass haloes and fewer merger events, we expect reduced Gini scores compared to a purely CDM universe. This is a promising result as a proof of concept. Future work will focus on a wider range of Milky Way-like haloes and galaxy histories to explore the significance of the variance of such statistics.
@@ 4.1 Simulations
Our simulations and subsequent dust post-processing show that we have potentially uncovered a new observational probe to test the nature of dark matter. Through the use of varying supernova efficiencies as our lever, we created simulations from different underlying dark matter types that produce the same “observable” satellite populations. These different supernova efficiencies lead to different dust properties in each halo. We find that the total dust mass produced by the fiducial CDM simulation, with a 10 per cent supernova efficiency, is 4 times greater than in the fiducial WDM simulation. We also presented a preliminary exploration of how dust is spatially distributed within the halo using the Gini statistic, which shows significant and persistent differences between dark matter models. The difference between the two suggests that dust can be used as an observational probe of the underlying dark matter. In reality, galactic dust masses are influenced by many factors, such as star formation histories (Calzetti et al., 2000; Clemens et al., 2013), mergers (Buck et al., 2023), and environments (Chelouche et al., 2007; Vogelsberger et al., 2019; Clark et al., 2023), so more work must be done before we draw conclusions about any observationally favoured model.
@@ 4.2 Milky Way satellites
There are many different sources for the mass of Milky Way satellites, which introduces some uncertainty into the correct values against which to calibrate the simulations. We use the masses published in Lovell & Zavala (2023) for all satellites included in the comparison in Fig. 1. This maintains a single source of data for simplicity and transparency, but does not explore the uncertainty of masses for these satellites. Many stellar masses are taken from McConnachie (2012), which has minor differences with other published masses for these satellites (Woo et al., 2008; Fattahi et al., 2018), though all are quite consistent for the larger satellites. Additionally, there is some scatter in the dynamical masses compared with Strigari et al. (2008) and Wolf et al. (2010), but once again, being the largest Milky Way satellites, the scatter in these measurements is relatively small.
Counter-intuitively, Fig. 1 shows that the fiducial WDM simulation produces more low-mass satellite galaxies than any other simulation, as mentioned in Section 3.2. This is rather unexpected, given that WDM is generally implemented to decrease satellite numbers. The apparent overproduction in the simulation 222Although we note that both WDM simulations produce fewer total subhaloes than their CDM counterparts. could be due to a number of factors, first among them the artificial fragmentation common in WDM simulations (Wang & White, 2007). We have explored removing such artificial haloes by implementing a minimum 100-particle limit. Our testing suggests that artificial haloes are not the cause of the excess. More work is needed to identify spurious haloes, e.g. identifying pancake haloes in the initial conditions (Lovell et al., 2014), but this is beyond the scope of this paper. Alternatively, Oman et al. (2024) found that WDM may boost the number of low-mass galaxies in 21-cm hydrogen observations. They state this may be because galaxies tend to form later in WDM simulations, and the subsequent delay in star formation helps them retain more gas than their CDM counterparts. Additionally, the higher concentration of CDM haloes may drive faster merging times, causing an excess of WDM satellites that have yet to fully merge by the same point in time. This may be true, at least for our Milky Way analogue. We defer such an investigation to later work, conscious that the most stringent limits on WDM particle masses are based on the observation of satellites below the resolution of these simulations (Banik et al., 2021; Nadler et al., 2021a; Dekker et al., 2022). However, the possibility of a WDM universe potentially producing more satellite galaxies in some haloes prompts further investigation into the variety of pathways for satellite accretion in a WDM universe and how these might compare with our own Milky Way satellites.